​Periodically, a town water department tests the the drinking water of homeowners for contaminants such as lead. The lead levels in water specimens collected for a sample of 10 residents of the town had a mean of 3.4​mg/L and a standard deviation of 2.1 ​mg/L. Complete parts a throughM c.

a.Construct a 90% confidence interval for the mean lead level in water specimens from the town.

​(Round to three decimal places as​ needed.)
Answer #1

Solution :

Given that,

Point estimate = sample mean = \bar x = 3.4

sample standard deviation = s = 2.1

sample size = n = 10

Degrees of freedom = df = n - 1 = 10 - 1 =9

At 90% confidence level the t is ,

\alpha = 1 - 90% = 1 - 0.90 = 0.10

\alpha / 2 = 0.10 / 2 = 0.05

t\alpha /2,df = t0.05,9 = 1.833

Margin of error = E = t\alpha/2,df * (s /\sqrtn)

= 1.833 * (2.1 / \sqrt 10)

= 1.217

The 90% confidence interval estimate of the population mean is,

\bar x - E < \mu < \bar x + E

< \mu <

3.4 - 1.217 < \mu < 3.4 + 1.217

(2.183 , 4.617)